# Lightening the Load: A Cluster-Based Framework for A Lower-Overhead, Provable Website Fingerprinting Defense

### v1

@@ VI-C Out-of-Training Evaluation
We split the closed-world dataset, which comprises 95 webpages, into two disjoint subsets of webpages. One subset (containing approximately half of the webpages) serves as the training set for the defense, while the other subset is reserved for the out-of-training evaluation, simulating previously unseen webpages.
@@
To compare the efficiency of different defense strategies under practical constraints, we fix a series of time overhead thresholds and, for each threshold, identify the configuration that yields the lowest bandwidth overhead while keeping the time overhead below the given limit. We report this value for three scenarios: (i) the original Tamaraw defense, (ii) Adaptive Tamaraw on out-of-training webpages, and (iii) Adaptive Tamaraw on in-training webpages. We threshold time overhead as it represents the maximum tolerance a user can be expected to accept for a defense: time delays directly affect user experience while bandwidth affects network load.
TABLE IV: Comparison of bandwidth overhead for Tamaraw and Adaptive Tamaraw under various fixed time overhead constraints. Results are reported for both in-training and out-of-training settings of Adaptive Tamaraw. Values in parentheses represent percentage reductions relative to the original Tamaraw. Adaptive Tamaraw consistently lowers both minimum and average bandwidth overhead, with the in-training variant achieving the largest reductions.
@@
Table IV presents the minimum bandwidth overhead achieved by each defense method under varying fixed time overhead thresholds. Adaptive Tamaraw on in-training webpages shows substantial gains over the original Tamaraw, especially under tight time constraints. For instance, when the time overhead is limited to 10%, Adaptive Tamaraw (in-training) achieves a 52% reduction in bandwidth overhead compared to the baseline. Notably, these improvements are more pronounced at lower time overhead thresholds, which is particularly practical: minimizing additional delay is often crucial for preserving user experience.
In the out-of-training setting, where all test traces are from previously unseen webpages, Adaptive Tamaraw still manages to slightly outperform the global Tamaraw strategy. The reductions are modest (e.g., 2–7%), but consistent across all thresholds. This suggests that even for webpages not present during training, partial trace similarity to known anonymity sets can still yield meaningful gains.
In addition to the minimum bandwidth values, we report the average bandwidth overhead, computed as the area under the curve (AUC) across all Pareto pairs, normalized by the range of time overheads. Both in-training and out-of-training variants of Adaptive Tamaraw demonstrate reduced average bandwidth overhead compared to the original Tamaraw baseline.
This experiment demonstrates that when Adaptive Tamaraw is trained on webpages the user would visit, it achieves its best performance, but it still outperforms Tamaraw even if it is not.
@@ VI-D1 Theoretical Bounds
In this section, we provide a numerical evaluation of the average success probability bound stated in Theorem C.2, which measures the level of security offered by Adaptive Tamaraw against any passive website fingerprinting adversary. We use the anonymity sets generated from the in-training evaluation (Section VI-B).
@@
Figure 6 presents the computed theoretical upper bounds on the attack success probability, as described in Theorem C.2, across varying anonymity set sizes $k$ and for three different values of the bucket length parameter $L\in\{100,500,1000\}$.
As expected, the attacker’s success probability decreases consistently with larger anonymity set sizes. This confirms that increasing $k$ enhances the uncertainty from the adversary’s perspective, thereby reducing the effectiveness of any passive WF attack. It is interesting to note that the decrease is not inverse to $k$; which says that larger $k$ encourages anonymity sets that have many traces from the same site. Additionally, increasing $L$ (padding traces to longer lengths) also suppresses maximum attacker accuracy. For instance, at $k=7$, the success probability drops from approximately 45% with $L=100$ to about 35% with $L=1000$, but note that the overhead cost is substantial as seen earlier in Table III.
Figure 6: Formal upper bound on an attacker’s accuracy for different anonymity-set sizes $k$ and bucket-length parameters $L$ in our Adaptive Tamaraw analysis (Section V-F). Increasing $k$ enlarges the pre-image of each defended trace, while a larger $L$ suppresses size-based leakage within each bucket. Together these levers tighten the bound from $\approx\!0.75$ at $(k,L)=(2,100)$ to below $0.17$ at $(k,L)=(30,1000)$, illustrating the trade-off between security and overhead.
We also evaluated the theoretical bound in the out-of-training scenario with $k=7$ and $L=100$. Specifically, we assigned each trace from the out-of-training dataset to its predicted anonymity set and computed the corresponding theoretical bound, then averaged the results over all traces. This yielded a maximum attacker accuracy bound of 31%, indicating that the theoretical framework remains valid even when test traces originate from webpages unseen during training. Notably, this bound is lower than its in-training counterpart in Figure 6 (45%), which can be attributed to the defense behavior in the out-of-training setting: as shown by the modest bandwidth improvements in Table IV, the defense tends to fall back toward behavior resembling the original, non-adaptive Tamaraw. As a result, the applied parameters are more uniform across traces, leading to more homogeneous traffic patterns and lower attack success rates.
@@ VI-D2 Empirical Attacks
In alignment with our deployment-oriented evaluation in Section VI-C, we use the practical configuration of $k=7$ and $L=100$. We randomly sample four $(\rho_{\text{in}},\rho_{\text{out}})$ pairs from the 33 Pareto-optimal global configurations and apply Adaptive Tamaraw to the training traces using each selected pair. We then conduct adversarial training using the defended traces and evaluate the attack success rate of each WF model on the corresponding test set. The training protocols and hyperparameters for all four models are the same mentioned in their original paper.
@@
TABLE V: Comparison of the theoretical maximum attacker accuracy bound with the empirical accuracy of various website fingerprinting attacks (kFP, Tik-Tok, RF, LASERBEAK) across four randomly selected $(\rho_{\text{out}},\rho_{\text{in}})$ padding rate pairs. Results are shown for $k=7$ and $L=100$.
@@ VII Discussion and Conclusion
We introduced a new defense framework that bridges regularization and supersequence methods with three core components: pattern-level clustering, $(k,l)$-diverse anonymity set construction, and early-time switching. We instantiated this design using Tamaraw, yielding Adaptive Tamaraw, a defense that maintains provable guarantees while adapting to trace-specific structure.
Based on our experiments, adaptive Tamaraw offers meaningful gains when tested on webpages it was trained on, and performs similarly to the original Tamaraw on webpages it was not trained on. As $k$ increases, the defense gradually resembles the classic Tamaraw, allowing control over the trade-off between privacy and overhead (Table  III and Figure 6).
While we instantiate our approach using Tamaraw due to its formal upper bound on information leakage, the core our approach is general and can be extended to other defenses that support fixed reugalarization configurations.
@@ Appendix A Model Architectures
Holmes. We adopt the Holmes model architecture proposed in [40] for early-stage website fingerprinting. Holmes utilizes a hybrid convolutional encoder trained via supervised contrastive learning to align partial traffic segments with their corresponding full-trace embeddings. The encoder consists of two initial 2D convolutional blocks, each composed of a Conv2D layer, batch normalization, and max pooling, followed by four 1D convolutional blocks comprising Conv1D layers with kernel size 3, batch normalization, and pooling operations. Residual connections are incorporated throughout the network to preserve spatial-temporal dependencies during downsampling. The model learns to embed partial traces such that their representations closely match those of complete traces from the same website, to achieve early prediction with limited input.
[…]
@@ B-A Tamaraw Parameter Grid and Pareto Frontier
To characterize the trade-off between bandwidth and time overhead in the Tamaraw defense, we constructed a grid of $(\rho_{\text{in}},\rho_{\text{out}})$ pairs following the method recommended in the original Tamaraw paper [11]. Specifically, we started with the original values used in that work:
$\rho_{\text{in}}^{\text{init}}=0.012,\quad\rho_{\text{out}}^{\text{init}}=0.04,$
and varied each parameter independently across 14 values, ranging from 7 times smaller to 7 times larger on a logarithmic scale. That is, each consecutive value differs from the previous by a multiplicative factor of $e^{\ln(7)/7}\approx 1.316$. This results in a total of $14\times 14=196$ combinations of $(\rho_{\text{in}},\rho_{\text{out}})$ pairs.
For each configuration, we applied the Tamaraw defense and measured both the bandwidth and time overhead using a batch of traces, fixing the parameter $L=100$. Among the resulting overhead pairs, we retained only the Pareto-optimal configurations (those for which no other parameter pair simultaneously achieves lower bandwidth and time overhead). The 33 obtained points form the Pareto frontier, representing the optimal trade-off curve between efficiency and latency.
Figure 7 illustrates the resulting Pareto frontier. Each red dot corresponds to a Pareto-optimal $(\rho_{\text{in}},\rho_{\text{out}})$ configuration, annotated with its corresponding (bandwidth, time) overhead pair. The curve reveals the inherent trade-off: significant reductions in one metric typically come at the cost of increases in the other.
@@
B-B Overhead Improvement Across Cluster Sizes
@@
To evaluate the effectiveness of adaptive switching, we measure the overhead reduction achieved by Adaptive Tamaraw compared to the static global Tamaraw defense. For each test trace, we first compute the total overhead (sum of bandwidth and time overhead) under both the global configuration and the adaptive configuration, and report the average difference across all traces. Figure 8 illustrates the average overhead improvement achieved by Adaptive Tamaraw over static global Tamaraw configurations, plotted as a function of the anonymity set size $k$. Results are shown for three different bucket-length parameters: $L\in\{100,500,1000\}$. The vertical axis reports the average total overhead reduction across all test traces and global configurations. Across all values of $L$, we observe that Adaptive Tamaraw consistently reduces overhead compared to the baseline. Smaller anonymity sets (i.e., decreasing $k$) generally lead to greater improvements, especially when $L$ is large. For $L=100$, the average overhead improvement steadily rises with $k$, reaching a reduction of approximately 15% at $k=30$. A similar but slightly more subdued trend is seen for $L=500$, where the maximum gain is around 25%. The most dramatic improvements are observed for $L=1000$, where average overhead savings exceed 60% at $k=30$ and remain significantly better than the baseline even at higher values of $k$.
This trend highlights the benefit of clustering traffic patterns into larger and more diverse anonymity sets, which allows the second phase of the defense to apply more relaxed parameters while preserving provable security. It also confirms that the adaptive switching strategy is particularly effective when the regularization budget (controlled by $L$) is high—leading to large opportunities for savings through early switching. Interestingly, the improvements are often most pronounced when $k$ is small (reaching around 100% reduction when $L=1000$, and 40% when $L=100$). This may be due to the fact that smaller anonymity sets contain more homogeneous traffic patterns, allowing the anonymity set–specific Tamaraw parameters to better optimize a narrower distribution of traces. As $k$ increases, the diversity within each set grows, which can dilute the effectiveness of a single shared configuration across that set.
@@
B-C Cluster Diversity Evaluation
As discussed in Section V-D, our proposed $k$-anonymity-based clustering algorithm relies on a Tamaraw-specific distance function that implicitly encourages low intra-cluster diversity. To empirically validate this behavior, we evaluate the purity of the anonymity sets generated by our clustering algorithm.
Following standard definitions [43], we compute the purity $p_{i}$ of the $i$-th cluster as:
$p_{i}\triangleq\frac{\max_{j}N_{i,j}}{N_{p_{i}}}\times 100\%,$
where $N_{i,j}$ is the number of traces in cluster $i$ that originate from website $j$, and $N_{p_{i}}$ is the total number of traces in cluster $i$. The purity thus reflects the dominance of the most frequent website in a given cluster, with lower values indicating greater diversity.
We compute the average purity across all clusters for each anonymity set size $k$ with fixed $L=100$. Figure 9 presents the resulting values alongside the theoretical lower bound $f(k)=1/k$, which represents the best achievable diversity for a perfectly balanced cluster.

### v2

@@ VI-C Out-of-Training Evaluation
We split the Sirinam et al. closed-world dataset, which comprises 95 webpages, into two disjoint subsets of webpages. One subset (containing approximately half of the webpages) serves as the training set for the defense, while the other subset is reserved for the out-of-training evaluation, simulating previously unseen webpages.
@@
To compare the efficiency of different defense strategies under practical user experience constraints, we evaluate their performance across a series of fixed time overhead constraints. For each constraint (e.g., requiring time overhead to be less than 10%), we identify the defense configuration from our Pareto-optimal set that yields the lowest possible bandwidth overhead while respecting that time limit. This methodology allows us to answer a practical question: Given a maximum acceptable delay, what is the minimum bandwidth cost? We report these values for three scenarios: (i) the original Tamaraw defense, (ii) Adaptive Tamaraw on out-of-training webpages, and (iii) Adaptive Tamaraw on in-training webpages.
Table IV presents the results of this analysis, where each column represents a different time overhead constraint. The values in the table show the minimum bandwidth overhead achieved by the best configuration that satisfies the given time constraint. Adaptive Tamaraw on in-training webpages shows substantial gains over the original Tamaraw, especially under tight time constraints. For instance, when the time overhead is limited to 10%, Adaptive Tamaraw (in-training) achieves a reduction of 52 percentage points in bandwidth overhead compared to the baseline. Notably, these improvements are more pronounced at lower time overhead thresholds, which is particularly practical: minimizing additional delay is often crucial for preserving user experience.
TABLE IV: Bandwidth overhead for Tamaraw and Adaptive Tamaraw under fixed time overhead constraints in both in-training and out-of-training settings. Parentheses show percentage reductions relative to Tamaraw. Adaptive Tamaraw consistently reduces minimum and average bandwidth overhead, with largest reductions in-training.
@@
In the out-of-training setting, where all test traces are from previously unseen webpages, Adaptive Tamaraw still manages to slightly outperform the global Tamaraw strategy. The reductions are modest (e.g., 2–7 percentage points), but consistent across all thresholds. This suggests that even for webpages not present during training, partial trace similarity to known anonymity sets can still yield meaningful gains.
Along with minimum bandwidth values, we report the average bandwidth overhead, computed as the normalized area under the Pareto curve. Both in-training and out-of-training versions of Adaptive Tamaraw show lower average overhead than the original Tamaraw. When Adaptive Tamaraw is trained on webpages the user would visit, it achieves its best performance, but it still outperforms Tamaraw even if it is not.
@@ VI-D1 Theoretical Bounds
In this section, we provide a numerical evaluation of the average success probability bound stated in Theorem E.2, which measures the level of security offered by Adaptive Tamaraw. We use the anonymity sets generated from the in-training evaluation (Section VI-B).
@@
Figure 6 presents the computed theoretical upper bounds on the attack success probability, as described in Theorem E.2, across varying anonymity set sizes $k$ and for three different values of the bucket length parameter $L\in\{100,500,1000\}$. As expected, the attacker’s success probability decreases with larger anonymity set sizes, confirming that increasing $k$ heightens adversarial uncertainty and weakens passive WF attacks. Interestingly, the drop is not inversely proportional to $k$, indicating that larger $k$ leads to sets containing many traces from the same site. Increasing $L$ (padding length) also lowers maximum attacker accuracy, for example, at $k=7$, success falls from about 45% with $L=100$ to 35% with $L=1000$, though with substantial overhead, as shown in Table III.
Figure 6: Formal upper bound on attacker accuracy for different anonymity-set sizes $k$ and bucket-length parameters $L$ (Section V-F). Increasing $k$ enlarges each trace’s pre-image, while larger $L$ suppresses size-based leakage. Together they tighten the bound from $\approx\!0.75$ at $(k,L)=(2,100)$ to below $0.17$ at $(k,L)=(30,1000)$, illustrating the security-overhead trade-off.
We also evaluated the theoretical bound in the out-of-training scenario with $k=7$ and $L=100$. We assigned each trace to its predicted anonymity set, computed the corresponding bound, and averaged across all traces. This yielded a maximum attacker accuracy of 31%, confirming the theoretical framework remains valid for unseen webpages. This bound is lower than the in-training counterpart in Figure 6 (45%), possibly because the defense reverts toward original Tamaraw behavior out-of-training, as evidenced by modest bandwidth improvements in Table IV. This produces more uniform parameters and homogeneous traffic patterns, lowering attack success rates.
@@ VI-D2 Empirical Attacks
In alignment with Section VI-C, we use the practical configuration of $k=7$ and $L=100$. We randomly sample four $(\rho_{\text{in}},\rho_{\text{out}})$ pairs from the 33 Pareto-optimal global configurations and apply Adaptive Tamaraw to the training traces using each selected pair. We then conduct adversarial training using the defended traces and evaluate the attack success rate of each WF model on the corresponding test set. The training protocols and hyperparameters for all four models are the same mentioned in their original paper.
@@
TABLE V: Theoretical maximum attacker accuracy bound versus empirical accuracy of website fingerprinting attacks for four randomly selected $(\rho_{\text{out}},\rho_{\text{in}})$ padding rate pairs with $k=7$ and $L=100$.
@@ VII Discussion and Conclusion
We proposed a defense framework combining regularization and supersequence methods with three components: pattern-level clustering, $(k,l)$-diverse anonymity sets, and early-time switching. Instantiated as Adaptive Tamaraw, it maintains provable guarantees while adapting to trace structure. Based on our experiments, adaptive Tamaraw offers meaningful gains when tested on webpages it was trained on, and performs similarly to the original Tamaraw on webpages it was not trained on. As $k$ increases, the defense approaches classic Tamaraw, enabling control over the privacy–overhead trade-off (Table III, Figure 6).
While we instantiate our approach using Tamaraw due to its formal security bound, our approach is general and can be extended to other defenses that support fixed regularization. Each stage of our approach presents opportunities for future research. For instance, more advanced representation learning could enhance pattern detection, while novel clustering algorithms could further optimize anonymity set generation. That said, the computational analysis in Appendix D shows that our current instantiation of Adaptive Tamaraw adds negligible latency and memory overhead on standard hardware. This makes it suitable for practical use in the Tor ecosystem, such as a Pluggable Transport in the WFDefProxy framework [15]. With inference latency under 2 ms, it can classify and adjust padding parameters in real time without computational bottlenecks, making it a viable option for Tor Browser integration.
@@ Appendix A Model Architectures
Holmes. We use the Holmes architecture from Deng et al. [10], a hybrid convolutional encoder with 2 blocks of Conv2D layers, followed by 4 blocks of Conv1D layers, trained with supervised contrastive learning to align partial traffic traces with full-trace embeddings for early-stage website fingerprinting.
k-Fingerprinting (kFP). We use kFP [16], a random forest with 100 trees per (website, safe time) pair, based on packet and timing statistics; fingerprints are leaf ID sequences classified by $k$-NN.
@@ B-A Tamaraw Parameter Grid and Pareto Frontier
To characterize the trade-off between bandwidth and time overhead in the Tamaraw defense, we constructed a grid of $(\rho_{\text{in}},\rho_{\text{out}})$ pairs following the method recommended in the original Tamaraw paper [5]. Specifically, we started with the original values used in that work: $\rho_{\text{in}}^{\text{init}}=0.012$ and $\rho_{\text{out}}^{\text{init}}=0.04$, and varied each parameter independently across 14 values, ranging from 7 times smaller to 7 times larger on a logarithmic scale. That is, each consecutive value differs from the previous by a multiplicative factor of $e^{\ln(7)/7}\approx 1.316$. This results in a total of $14\times 14=196$ combinations of $(\rho_{\text{in}},\rho_{\text{out}})$ pairs.
For each configuration, we applied Tamaraw and measured the overheads on the traces. We kept only Pareto-optimal pairs, those where no other configuration achieves lower overhead in both metrics. Figure 7 shows the 33 points fpr Sirinam et al. frontiers for $L=100$. Each red dot marks a Pareto-optimal $(\rho_{\text{in}},\rho_{\text{out}})$ with its (bandwidth, time) overhead.
@@
B-B Security Bound Analysis
Our security bound in Section V-F is designed to bound the attacker’s average success rate, assuming the attacker fully knows the trace and the underlying trace information, including the switching time $\tau_{s}$, the chosen anonymity set and the sequence length. Here we evaluate how much information is individually leaked by a $\tau_{s}$-aware adversary who, based on the observed switching time and post-switch rate pair, can infer the corresponding anonymity set. We performed this evaluation across all 33 Pareto-optimal global configurations (with $L=100$ and $k=7$), comparing the empirical success rate of this adversary against the theoretical security bound derived in Section V-F. All defense parameters, including anonymity sets and safe timestamps, are derived from the training set, and the adversary’s performance is measured on the held-out test set.
Our analysis reveals that the switching mechanism does indeed leak the anonymity set identity with high probability: on average, the correct anonymity set accounts for 95.64% of the traces sharing a specific $(\tau_{s},\text{rates})$ tuple. Our theoretical security bound has already assumed such leakage. As shown in Figure 8, the $\tau$-aware attacker’s empirical accuracy consistently remains below the theoretical security bound across all configurations. This confirms that even when the adversary successfully identifies the anonymity set via the switching signal, the $k$-anonymity and $l$-diversity properties within that set remain sufficient to cap their success rate. Furthermore, a gap exists between the theoretical bound and the attacker’s actual performance, which might also suggest that the presence of misclassified and non-switching traces introduces additional ambiguity, further confusing the adversary.
Figure 8: Comparison of the theoretical security bound and the empirical accuracy of a $\tau$-aware attacker across 33 global configurations ($L=100,k=7$). The attacker consistently performs below the bound, demonstrating that the defense remains robust even when the switching signal reveals the anonymity set.
B-C Generalization to Onion Service Traffic
To further evaluate the robustness of our method in out-of-training conditions, we conduct an experiment using traces from onion services, which can exhibit different traffic characteristics than standard websites. We use the onion service dataset collected by Overdorf et al. [27], which contains traffic from distinct services, from which we use 100 sites. The experimental setup is the same as our primary out-of-training evaluation: we construct anonymity sets using only the Sirinam et al. dataset, and then apply both the static and adaptive defenses to the unseen onion traces to assess our method’s ability to generalize to this entirely different domain. Table VI compares the minimum bandwidth overhead between Adaptive Tamaraw and static Tamaraw under fixed time overhead constraints. For ease of comparison, we have also included the results for normal out-of-training websites from Section VI.C. The data shows that Adaptive Tamaraw provides improvements in both scenarios. For normal out-of-training websites, the adaptive approach yields modest but consistent reductions in bandwidth overhead of up to 5 percentage points. Similarly, Adaptive Tamaraw reduces bandwidth overhead across all constraints on the onion service traffic. Under a 20% time overhead limit, our approach reduces bandwidth overhead by 8 percentage points for onion websites, demonstrating that even with fundamentally different website traffic, our adaptive mechanism identifies structural similarities to provide improvements.
TABLE VI: Comparison of minimum bandwidth overhead for Tamaraw and Adaptive Tamaraw under various fixed time overhead constraints when applied on out-of-training websites (both normal and onion websites). Values in parentheses represent percentage point reductions relative to Tamaraw.
Website Method Minimum Bandwidth Overhead (%)
<<20% <<45% <<125% <<200%
Normal Tamaraw 168 124 83 67
Adaptive 166 (-2) 119 (-5) 80 (-3) 67 (0)
Onion Tamaraw 195 120 78 66
Adaptive 187 (-8) 116 (-4) 77 (-1) 65 (-1)
B-D Overhead Improvement Across Cluster Sizes
@@
We present the overhead reduction achieved by Adaptive Tamaraw compared to Tamaraw depending on cluster sizes. For each test trace in the Sirinam et al. dataset, we compute the average improvement for bandwidth and time overhead individually and average the results. Figure 9 illustrates the average overhead improvement achieved by Adaptive Tamaraw over static global Tamaraw configurations, plotted as a function of the cluster size $k$. Across all values of $L$, we observe that Adaptive Tamaraw consistently reduces overhead compared to the baseline. The largest reductions (observed for $L=1000$) are seen in the time overhead, which is cut by over 50 percentage points for small $k$. For smaller bucket lengths ($L=100$ and $L=500$), the savings are more modest but are more evenly distributed between both bandwidth and time improvements. In all scenarios, smaller anonymity sets (i.e., decreasing $k$) generally lead to greater improvements, and furthermore when $L$ is large.
@@
Appendix C Ablation Studies
C-A Cluster Diversity Evaluation
As discussed in Section V-D, our $k$-anonymity clustering algorithm uses a Tamaraw-specific distance function that encourages low intra-cluster diversity. To validate this, we evaluate the purity of generated anonymity sets.
Following [35], we compute cluster $i$’s purity $p_{i}$ as $\frac{\max_{j}N_{i,j}}{N_{p_{i}}}$, where $N_{i,j}$ is the number of traces from website $j$ in cluster $i$, and $N_{p_{i}}$ is the total traces in cluster $i$. Lower purity indicates greater diversity. We compute average purity across all clusters for each $k$ with fixed $L=100$. Figure 10 shows these values alongside the theoretical lower bound $f(k)=1/k$, representing optimal diversity for a perfectly balanced cluster. As shown in the figure, the empirical purity values closely track the $1/k$ curve, indicating that our clustering procedure approaches the theoretical optimum in terms of diversity.